OpenStudy (anonymous):

calculate CONSUMERS Surplus: q=50-3p; p(bar)=6 the equation is: CS=(the integration from 0 to q(bar))=[D(q)-p(bar)] dq PLEASE PLEASE PLEASE help.

OpenStudy (anonymous):

\[\int\limits_{0}^{32}(32-q)/3 dq\]

OpenStudy (anonymous):

\[q=50-3p\] \[\bar{p}=6\] \[\int\limits_{0}^{\bar{q}} Dq-\bar{p} \ dq\]

OpenStudy (anonymous):

is this it?

OpenStudy (anonymous):

you need to put the q=50-3p into terms of p not q. how do I do that?

OpenStudy (anonymous):

so just solve for p q=50-3p subtract 50 from both sides then divide by -3 to both sides then you should get p in terms of q